Determine the amount of an investment if $15,000 is invested at an iterest of 4.85% compound quarterly for 5 years
Found 2 solutions by checkley77, nyc_function:
Answer by checkley77(12844)

15,000(1+.0485/4)^5*4
15,000(1+.012125)^20
15,000*1.012125^20
15,000*1.2726=19,088.61 ans.

Answer by nyc_function(2741)

Use the formula I = P(1 + r/4)^4, where P = 15,000 and r = 0.0485.
Plug into formula and simplify using your calculator.
We are searching for I or interest.
I = 15,000(1 + 0.0485/4)^4
Can you do it now?
NOTE: The number 4 represents interest every 4 months since this is a quarterly compounding question.